# Agentic Program Verification

### v1

@@ 6. Experimental Results
6.1. RQ.1: Effectiveness of AutoRocq
@@
For this research question, we investigate AutoRocq’s effectiveness in proof generation on both (a) a non-program-related subset of lemmas from CoqGym, and (b) new lemmas extracted from SV-COMP programs. We report the number of proved lemmas from each tool in 4(b). Figure LABEL:fig:rq1_succ shows the total number of lemmas successfully proved by each tool. Figure LABEL:fig:rq1_venn presents the detailed breakdown of the lemmas proven by each tool, on CoqGym (left) and SV-COMP programs (right). We note that PALM’s results (in green) may not reflect its full capability, as several lemmas in the benchmarks are incompatible with Rocq 8.12.
@@
Lemmas from CoqGym. From Figure LABEL:fig:rq1_succ, we observe that AutoRocq successfully proves 319 of them (51.1%), outperforming other tools by a large margin of 20.8% (compared with Rango (rango-icse25)) to 343.0% (compared with PALM (lu2025palm)). Among the baseline tools, Rango  performs the best on this dataset, synthesizing proofs for 264 (42.3%) of the lemmas on CoqGym. The Venn diagram (Figure LABEL:fig:rq1_venn, left) also demonstrates that AutoRocq proves the most lemmas (98) uniquely. Our tool is followed by Rango and PALM, which account for 30 and 18 uniquely proved lemmas, respectively. The results suggest that AutoRocq is highly effective in automatic proof generation.
Lemmas from SV-COMP programs. More importantly, we study complex proof obligations that represent real verification workflows. These lemmas from SV-COMP programs indeed pose a greater challenge to all LLM-based tools, namely AutoRocq, Rango, and PALM, as evidenced by their lower success rates overall. AutoRocq manages to prove 198 lemmas (30.9%) from the benchmark, which is significantly higher than all other baseline tools. Results produced by Rango, QEDC, and P9001 lie in close proximity with one another, with each producing proofs for around 135 lemmas. AutoRocq outperforms them by 42.4%, 47.8%, and 51.1%, respectively. The number of lemmas proved by AutoRocq is also 204.6% higher than that of PALM.
Figure LABEL:fig:rq1_venn (right) shows that AutoRocq generates proof scripts for 44 unique lemmas (owing to its agentic workflow) that no other tool can prove. In contrast, no other tool is able to generate proofs uniquely for more than 6 lemmas. In fact, only 23 lemmas proved by any other baselines elude AutoRocq! We note that there is a large overlap among the successes of baseline tools. There are 106 lemmas that are proved commonly by at least three of the baselines, suggesting a strong convergence in capabilities among them.
(b) [RQ.1] Lemmas proved by each tool on CoqGym and SV-COMP programs. On both benchmarks, AutoRocq is able to prove more lemmas, and has the most number of uniquely proved lemmas.
@@
(c) [RQ.1] # of lemmas proved from SV-COMP programs: a breakdown by lemmas’ complexity.
@@
To better understand the high effectiveness of AutoRocq, we further correlate the number of successfully proved SV-COMP lemmas with their complexity. 4(c) details the breakdown of successful attempts from different tools, grouped into five complexity buckets. We present the breakdown for both the term count (left) and the hypothesis count (right) in the original goal. Buckets on the right correspond to lemmas containing more terms/hypotheses, and thus are typically more structurally and contextually complicated. Results reveal an emerging pattern: as the original goal becomes more verbose and richer in context, the number of successfully proved lemmas decreases in general. However, AutoRocq retains its relative effectiveness in proving longer, more complex goals, resulting in a uniform lead across different complexity levels.
@@
Additionally, we report the breakdown by the categories of lemmas, as shown in Figure 4. Results show that AutoRocq is particularly effective in proving non-overflow-related lemmas and loop invariants in the program, outperforming the best baselines by 53% and 22%, respectively. On lemmas related to functional correctness, AutoRocq also performs comparably with three baseline approaches. Results indicate AutoRocq’s performance is consistent for verifying different types of program properties, showing its versatility and generality.
@@
We also compare the efficiency of different tools in terms of the time taken to generate a successful proof on the SV-COMP lemmas. On average, AutoRocq takes 21.3 seconds to generate a successful proof, expending less time than other LLM-based approaches, namely Rango (105.5 seconds) and PALM (45.4 seconds). AutoRocq’s result is also comparable to P9001 (22.6 seconds). Notably, QEDC finishes extremely fast (5.1 seconds) when it does succeed, thanks to its well-optimized tactic-prediction model for the tactic generation.
@@
Answer to RQ.1: Overall, AutoRocq is significantly more effective in proving obligations extracted from real C programs (SV-COMP), outperforming baselines by 42%–205%. Notably, AutoRocq proves 44 lemmas that none of the other approaches can prove. It also outperforms the baselines by a similar margin on mathematical lemmas from CoqGym.
Figure 4. [RQ.1] # of lemmas proved from SV-COMP programs: a breakdown by the category of lemmas.
Figure 5. [RQ.2] The log frequency of invocation for context search commands on SV-COMP programs.
6.2. RQ.2: Design Choices of AutoRocq
To better understand the interplay between AutoRocq’s components, we evaluate how different design choices affect the overall effectiveness by implementing 6 variants of AutoRocq. We conduct the evaluation on 70 randomly sampled lemmas from the SV-COMP programs. Specifically, we examine the following variant approaches:
@@
  • $\lnot EF$}, $\lnot HM$}: No error feedback or history manager (Section 4.3).
@@
We report the success rates, and AutoRocq’s relative improvement, of different variants in Table 2. Results show that all variants only prove a fraction of lemmas compared to full-fledged AutoRocq. In particular, AutoRocq sees the most significant improvement from the ablative version $\lnot CS$}, which replaces the context search component with a direct request of tactic through prompting. This indicates that agentic access to the proof context contributes the most to our approach’s effectiveness. Intriguingly, querying for contexts much more frequently ($Err_{1}$}) or less frequently ($Err_{5}$}) also affects the results negatively, suggesting that the amount of contexts supplied to the agent is paramount in practice. Removing other components in our approach, namely the proof-tree representation ($\lnot PT$}) or the feedback mechanisms ($\lnot EF$} and $\lnot HM$}), each reduces the efficacy moderately by $\sim 10\%-15\%$. We conclude that all components of AutoRocq are useful.
Table 2. [RQ.2] Success rates of different variants of AutoRocq on verifying sampled SV-COMP programs. All components contribute to the overall effectiveness, with context search ($\lnot CS$}) being the most significant.
@@
Answer to RQ.2: Each component of AutoRocq contributes to its overall effectiveness. AutoRocq’s agentic context search enhances its efficacy most significantly.
6.3. RQ.3: Comparison to Human-written Proofs
Since there is no ground truth for the lemmas extracted from SV-COMP programs, we cannot directly compare the proofs synthesized by AutoRocq with human-written proofs. In this subsection, we closely examine a proof synthesized by AutoRocq, and compare it to a manually crafted proof by a Rocq expert. The expert has more than six years of formal verification experience and one year of practical experience with Rocq. We use the same theorem wp_goal as presented in Figure 1(a), and report the human-written proof in 7(a).
To construct the proof in Figure LABEL:fig:proof-human, the expert had first to recognize that the problem’s core reduces to the mathematical principle that, for any integer $a$, $a^{2}\geq 100$ implies $|a|\geq 10$. This insight is non-trivial, as it requires moving beyond a direct case-based enumeration of values to grasp the underlying structural inequality. Only after achieving this conceptual leap could the expert elegantly bifurcate the problem based on the sign of i1 and construct the two custom, symmetric helper lemmas (lines 9–15 in Figure LABEL:fig:theorem) to complete the solution. Crafting this proof takes the Rocq expert 20 minutes of time.
(a) [RQ.3] Proof for wp_goal in Figure 1(a), written by a human expert prover in 20 minutes.
@@
Answer to RQ.3: With the help of context queries, AutoRocq is able to prove challenging lemmas, sometimes in even fewer steps than human experts. It also takes much less time.
@@ 6.4. Threats to Validity
LLMs may give incomplete or factually wrong responses to questions. As a result, verifying programs with untrusted LLMs may seem off-putting at first glance. However, we note that the certificate essentially comes from the trusted kernel of Rocq, which ensures that only derivations conforming to the formal rules of the proof assistant are accepted (coq-kernel). Therefore, all the generated proof scripts are true positives and are trustworthy, as any erroneous or misleading outputs from the LLMs will simply fail to be verified. Throughout our experiments, it takes \$1.22 on average to prove a single lemma without token caching. We deem this cost acceptable, compared to the costly human involvement in lemma-proving activities.
@@ 7. Case Study: Verifying Linux Kernel Modules
Table 3. [Cast Study] Verifying Linux kernel modules: # of lemmas proved, and the average time/steps (# of tactics) expended on proved lemmas. Across all tools, proofs are constructed for 13 unique lemmas. AutoRocq proves 12 (92%) of them, or 20% of all 60 lemmas.
Tools AutoRocq Rango PALM QEDC P9001 Combined
Proved? (max. 60) 12 3 10 3 2 13
Avg. Time (sec) 29.9 96.2 39.4 2.0 160.8
Avg. Steps 8.3 32.0 32.0 979.0
AutoRocq has demonstrated remarkable efficacy in automatically synthesizing rigorous proofs for lemmas related to smaller programs or a mathematical context. In this case study, we assess if AutoRocq can handle the complexity of real-world software. To this end, we build on earlier efforts (volkov2018lemma; efremov2018deductive) to reason about properties in the Linux kernel. Specifically, we examine 60 lemmas that formalize functional correctness in Linux kernel code. These lemmas reason about the correctness of intricate program behaviors, and come from various source files, such as memory management (e.g. memmove) and utilities such as string operations (e.g. strcpy) and type conversion (e.g. hex2bin). These lemmas tend to be significantly more involved and may contain up to hundreds of terms.
We report the statistics in Table 3. Results show that AutoRocq is the most effective, being able to synthesize the proofs for 12/60 (20%) lemmas. AutoRocq fails to construct a proof for only one lemma that is proved by any other tools (in this case PALM and P9001). In contrast, Rango, QEDC, and P9001 perform poorly on these tasks, which only manage to succeed in 3, 3, 2 lemmas, respectively. We also note that QEDC runs exceptionally fast when it is able to generate a proof, being 15x faster than AutoRocq, which is the second fastest. This suggests that QEDC’s search heuristic performs extremely well on certain cases. Nonetheless, such cases seem rare when verifying real, complex software, as it only succeeds on 3 lemmas.
Notably, PALM also achieves remarkable effectiveness and manages to prove 10/60 lemmas. Upon manual examination, we note that all successful proofs generated by PALM rely on heavy use of CoqHammer (coqhammer), a plug-in that directly discharges remaining subgoals to automated theorem provers (ATPs). CoqHammer is invoked by PALM as a last resort when none of its deterministic tactics repairs or resolves the encountered errors. AutoRocq currently does not rely on any external ATP to boost its performance. We note that more tools, such as with ATPs, could be easily incorporated into our agentic workflow, which we leave as future work.
AutoRocq can be applied to verify properties in complex software such as the Linux kernel. Access to more tools, e.g., automated theorem provers, may further improve our proof agent.

### v3

@@ 6. Experimental Results
6.1. RQ.1: Efficacy of AutoRocq on Mathematical Lemmas
@@
For this research question, we investigate AutoRocq’s effectiveness in proof generation on mathematical lemmas. We report the number of proved lemmas from each tool and a detailed breakdown in Figure 4. The bar plot (left) shows the total number of lemmas successfully proved by each tool, whereas the Venn diagram (right) presents the detailed breakdown of the proved lemmas. We note that PALM’s results may not reflect its full capability, as several lemmas in the benchmarks are incompatible with Rocq 8.12.
@@
The bar plot (Figure 4, left) shows that, across all 1,717 lemmas, AutoRocq successfully proves 824 (48.0%) of them, outperforming other tools by a large margin of 15.2% (compared with Rango) to 172.8% (compared with QEDC). Among the baseline tools, Rango performs the best on this dataset, synthesizing proofs for 715 (41.6%) of the lemmas on CoqGym. The two baselines using GPT-4.1 as the backend model, PALM and Copra, can both prove more than 500 mathematical lemmas. The Venn diagram (Figure 4, right) also demonstrates that AutoRocq proves the most lemmas (140) uniquely. Our tool is followed by PALM and Rango, which account for 78 and 38 uniquely proved lemmas, respectively. Copra is able to find proofs for 19 lemmas uniquely. The results suggest that AutoRocq is highly effective in automatic proof generation.
Figure 4. [RQ.1] Mathematical lemmas from CoqGym proved by each tool and their breakdown.
@@
Figure 5. [RQ.2] Verification lemmas from SV-COMP programs proved by each tool and their breakdown.
6.2. RQ.2: Efficacy of AutoRocq on Verification Lemmas
More importantly, we study complex proof obligations that represent real verification workflows, and present the results in Figure 5, with the bar plot (left) and its breakdown in the form of a Venn diagram (right). Overall, AutoRocq manages to prove 198 lemmas (30.9%) from the benchmark, outperforming the best baseline, Copra, by 10.6%. Results produced by Rango, QEDC, and P9001 lie in close proximity to one another, with each producing proofs for around 135 lemmas. AutoRocq outperforms them by 42.4%, 47.8%, and 51.1%, respectively. The number of lemmas proved by AutoRocq is also 204.6% higher than that of PALM.
The Venn diagram (Figure 5, right) shows that AutoRocq generates proof scripts for 28 unique lemmas (owing to its agentic workflow) that no other tool can prove. In contrast, no other tool is able to generate proofs uniquely for more than 6 lemmas. In fact, only 19 lemmas proved by any other baselines elude AutoRocq! We note that there is a large overlap among the successes of baseline tools. 113 lemmas are proved commonly by at least three of the baselines, suggesting a strong convergence in capabilities among them.
Compared to the results in Figure 4, the verification lemmas indeed pose a greater challenge to all neural theorem provers (including AutoRocq), as evidenced by their lower success rates overall. The relative degradation in performance is the most significant for PALM, indicating that whole-proof generation scales poorly as the lemmas grow in complexity. Similarly, Rango is the best-performing baseline on mathematical lemmas, but is outperformed by Copra in the SV-COMP benchmark. This performance decline can be attributed to a distribution mismatch: its underlying model was fine-tuned primarily on CoqStoq, and consequently struggles when encountering the more complex verification lemmas that lie outside this training data. In both benchmarks, AutoRocq proves the most lemmas and has the most unique successes, strongly suggesting its efficacy and robustness in automated proof generation.
Answer to RQ.1&2: Overall, AutoRocq is more effective in proving both mathematical and verification-derived lemmas, outperforming baselines by 15.2% to 172.8% on CoqGym and 10.6%–204.6% on SV-COMP. Notably, AutoRocq proves 168 lemmas (140 on CoqGym and 28 on SV-COMP) that none of the other approaches can prove.
@@
To better understand the remarkable efficacy of AutoRocq, we further correlate the number of successfully proved SV-COMP lemmas with their complexity. Figure 6 details the breakdown of successful attempts from different tools, grouped into five complexity buckets. We present the breakdown for both the term count (Figure 6(a)) and the hypothesis count (Figure 6(b)) in the original goal. Buckets on the right correspond to lemmas containing more terms/hypotheses, and thus are more structurally and contextually complicated. Results reveal an emerging pattern: as the original goal becomes more verbose and context-rich, the number of successfully proved lemmas generally decreases. However, AutoRocq retains its relative effectiveness in proving longer, more complex goals, resulting in an almost uniform lead across different complexity levels. Among the baselines, Copra scales with increased complexity better, and even manages to prove slightly more lemmas than AutoRocq on moderately difficult lemmas (150–224 terms or 15–22 hypotheses). This further demonstrates the advantage of feedback-guided proof generation in program verification.
(a) Breakdown by term count.
(b) Breakdown by hypothesis count.
Figure 6. [RQ.2] # of verification lemmas proved from SV-COMP programs: a breakdown by lemmas’ complexity.
@@
Additionally, we report the breakdown by the categories of lemmas, as shown in Figure 7. Results show that AutoRocq is particularly effective in proving non-overflow-related lemmas in the program, outperforming the best baselines by 37%. On lemmas related to loop invariants, AutoRocq is comparable with Copra and outperforms other baselines; whereas on lemmas specifying functional correctness, all approaches except PALM perform comparably. Results indicate that AutoRocq’s performance is consistent across different properties, demonstrating its versatility and generality.
@@
We also compare the efficiency of different tools in terms of the time taken to generate a successful proof on the SV-COMP lemmas. On average, AutoRocq takes 21.3 seconds to generate a successful proof, expending less time than other LLM-based approaches, namely Rango (105.5 seconds), PALM (45.4 seconds), and Copra (49.1 seconds). AutoRocq’s result is also comparable to P9001 (22.6 seconds). Notably, QEDC finishes extremely fast (5.1 seconds) when it does succeed, thanks to its well-optimized tactic-prediction model for the tactic generation.
@@
Figure 7. [RQ.2] # of lemmas proved from SV-COMP programs: a breakdown by the category of lemmas.
Figure 8. [RQ.3] The log frequency of invocation for context search commands on SV-COMP programs.
6.3. RQ.3: Design Choices of AutoRocq
To better understand the interplay between AutoRocq’s components, we evaluate how different design choices affect the overall effectiveness by implementing 6 variants of AutoRocq. We conduct the evaluation on 70 randomly sampled verification lemmas from SV-COMP programs. Specifically, we examine the following variant approaches:
@@
  • $\lnot EF$}, $\lnot HM$}: No error feedback or history manager for the on-the-fly learning (Section 4.3).
@@
We report the success rates and AutoRocq’s relative improvement, of different variants in Table 3. Results show that all variants only prove a fraction of lemmas compared to full-fledged AutoRocq. In particular, AutoRocq sees the most significant improvement from the ablative version $\lnot CS$}, which replaces the context search component with a direct request of tactic through prompting. This indicates that agentic access to the proof context contributes the most to our approach’s effectiveness. Intriguingly, querying for contexts much more frequently ($Err_{1}$}) or less frequently ($Err_{5}$}) also affects the results negatively, suggesting that the amount of contexts supplied to the agent is paramount in practice. Removing other components in our approach, namely the proof-tree representation ($\lnot PT$}) or the feedback mechanisms ($\lnot EF$} and $\lnot HM$}), each reduces the efficacy moderately by $\sim 10\%-15\%$. We conclude that all components of AutoRocq are useful.
Answer to RQ.3: Each component of AutoRocq contributes to its overall effectiveness. AutoRocq’s agentic context search enhances its efficacy most significantly.
Table 3. [RQ.3] Success rates of different variants of AutoRocq on proving sampled verification lemmas.
@@
6.4. RQ.4: Comparison to Human-written Proofs
Since there is no ground truth for the lemmas extracted from SV-COMP programs, we cannot directly compare the proofs synthesized by AutoRocq with human-written proofs. In this subsection, we closely examine a proof synthesized by AutoRocq, and compare it to a manually crafted proof by a Rocq expert. The expert has more than six years of formal verification experience and one year of practical experience with Rocq. We use the same theorem wp_goal as presented in Figure 1(a), and report the human-written proof in Figure 9.
To construct the proof in Figure 9(b), the expert had first to recognize that the problem’s core reduces to the mathematical principle that, for any integer $a$, $a^{2}\geq 100$ implies $|a|\geq 10$. This insight is non-trivial, as it requires moving beyond a direct case-based enumeration of values to grasp the underlying structural inequality. Only after achieving this conceptual leap could the expert elegantly bifurcate the problem based on the sign of i1 and construct the two custom, symmetric helper lemmas (lines 9–15 in Figure 9(a)) to complete the solution. Crafting this proof takes the Rocq expert 20 minutes of time.
@@
1Theorem wp_goal :
2 forall (i:Numbers.BinNums.Z) (i1:Numbers.BinNums.Z),
3 let x := (i1 * i1)%Z in ((-9%Z)%Z <= i1)%Z $\rightarrow\;$
4 ((-9%Z)%Z <= i)%Z $\rightarrow\;$ (i <= 9%Z)%Z $\rightarrow\;$ (i1 <= 10%Z)%Z $\rightarrow\;$
5 (i <= 10%Z)%Z $\rightarrow\;$((-2147483648%Z)%Z <= x)%Z $\rightarrow\;$
6 (100%Z <= x)%Z $\rightarrow\;$ (x <= 2147483647%Z)%Z $\rightarrow\;$ is_sint32 i1 $\rightarrow\;$
7 is_sint32 i $\rightarrow\;$ (i1 = 10%Z).
8
9(* Helper lemmas written by the human prover*)
10Lemma square_le_mono_nonneg :
11 forall a b, 0 <= a $\rightarrow\;$ 0 <= b $\rightarrow\;$ a * a <= b * b $\rightarrow\;$ a <= b.
12Proof. intros a b Ha Hb Hsq. nia. Qed.
13Lemma square_le_mono_neg :
14 forall a b, a < 0 $\rightarrow\;$ b < 0 $\rightarrow\;$ a * a <= b * b $\rightarrow\;$ b <= a.
15Proof. intros a b Ha Hb Hsq. nia. Qed.
(a) wp_goal in Figure 1(a) with human helper lemmas.
1Proof.
2 intros. subst x. assert (Hcases: 0 <= i1 \/ i1 < 0) by lia.
3 destruct Hcases as [Hge0 | Hlt0].
4 - assert (H10: 0 <= 10) by lia.
5 pose square_le_mono_nonneg.
6 specialize (l 10 i1 H10 Hge0).
7 simpl in l.
8 specialize (l H5). lia.
9 - pose square_le_mono_neg.
10 specialize (l (-10) i1).
11 assert (Hn10: -10 < 0) by lia.
12 specialize (l Hn10 Hlt0).
13 simpl in l.
14 specialize (l H5). lia.
15Qed.
(b) Main proof script written by a human Rocq expert.
Figure 9. [RQ.4] Proof for wp_goal in Figure 1(a), written by a human expert prover in 20 minutes.
Answer to RQ.4: With the help of context queries, AutoRocq is able to prove challenging lemmas, sometimes in even fewer steps than human experts. It also takes much less time.
@@ 6.5. Threats to Validity
LLMs may give incomplete or factually wrong responses to questions. As a result, verifying programs with untrusted LLMs may seem off-putting at first glance. However, we note that the certificate essentially comes from the trusted kernel of Rocq, which ensures that only derivations conforming to the formal rules of the proof assistant are accepted (Sozeau et al., 2025). Therefore, all the generated proof scripts are true positives and are trustworthy, as any erroneous or misleading outputs from the LLMs will simply fail to be verified. Throughout our experiments, it takes \$1.22 on average to prove a single lemma without token caching. We deem this cost acceptable, compared to the costly human involvement in lemma-proving activities.
@@
Impact of LLMs Used. Since AutoRocq adopts a general-purpose LLM (i.e., GPT-4.1) while some baseline approaches adopt custom ones, we want to study the impact of the underlying model. We conducted an additional comparison between AutoRocq and a variant of Rango (the best-performing baseline with a custom model) named Rango-GPT-4.1. In this variant, the fine-tuned LLM was replaced with GPT-4.1. Experimental results show that Rango-GPT-4.1 can only prove 11.43% of the verification lemmas used in Section 6.3, whereas AutoRocq proves 44.29%. These results strongly indicate that AutoRocq’s relative efficacy over competing tools comes from the agentic layer rather than from a powerful choice of LLM.
Impact of Rocq Versions. During evaluation, we used a different Rocq 8.12 for PALM, which is different from other comparative tools (i.e., Rocq 8.18), raising concerns about how different Rocq versions may affect the experimental results. In theory, Rocq version differences do not materially affect the comparison between the PALM and AutoRocq, as most of the version changes concern syntax, parsing, or library organization, rather than the proof engine that our method relies on. Practically, the impact would be minimal, as AutoRocq dynamically retrieves and queries lemmas from the current Rocq environment and thus does not depend on version-specific assumptions.
@@ 7. Case Study: Verifying Linux Kernel Modules
Table 4. [Cast Study] Verifying functional correctness in Linux kernel modules: # of lemmas proved, and the average time (in seconds) and steps (in # of tactics) expended on proved lemmas.
Tools AutoRocq Rango PALM Copra QEDC P9001 AutoRocq-w/H
Proved? (max. 60) 12 3 10 7 3 2 18
Avg. Time 29.9 96.2 39.4 65.1 2.0 160.8 65.9
Avg. Steps 14.3 32.0 50.0 32.0 979.0 16.7
AutoRocq has demonstrated remarkable efficacy in automatically synthesizing rigorous proofs for lemmas related to smaller programs or a mathematical context. In this case study, we assess if AutoRocq can handle the complexity of real-world software. To this end, we build on earlier efforts (Volkov et al., 2018; Efremov et al., 2018) to reason about properties in the Linux kernel. Specifically, we examine 60 lemmas that formalize functional correctness in Linux kernel code. These lemmas reason about the correctness of intricate program behaviors, and come from various source files, such as memory management (e.g., memmove) and utilities such as string operations (e.g., strcpy) and type conversion (e.g., hex2bin). These lemmas tend to be significantly more involved and may contain up to hundreds of terms.
We report the statistics in Table 4. Results show that AutoRocq is the most effective, being able to synthesize the proofs for 12/60 (20%) lemmas. Copra manages to prove 7 lemmas as well. In contrast, Rango, QEDC, and P9001 perform poorly on these tasks, which only manage to succeed in 3, 3, 2 lemmas, respectively. We also note that QEDC runs exceptionally fast when it is able to generate a proof, being 15x faster than AutoRocq, which is the second fastest. This suggests that QEDC’s search heuristic performs extremely well on certain cases. Nonetheless, such cases seem rare when verifying real, complex software, as it only succeeds on 3 lemmas.
Notably, PALM also achieves remarkable effectiveness and proves 10/60 lemmas. Upon manual examination, we note that all successful proofs generated by PALM rely on heavy use of CoqHammer (Czajka and Kaliszyk, 2018), a plug-in that directly discharges remaining subgoals to automated theorem provers (ATPs). CoqHammer is invoked by PALM as a last resort when none of its deterministic tactics repairs or resolves the encountered errors. Based on this finding, we create a new variant, AutoRocq-w/H, by augmenting our approach with basic CoqHammer integration. Specifically, the hammer is invoked with its default settings once per proof state upon the first failed tactic application. Such integration yields significant efficacy gains, with 18 lemmas (30%) proved, indicating that access to additional tools could further enhance the capabilities of our proof agent.
AutoRocq can be applied to verify properties in complex software such as the Linux kernel. Access to external automated theorem provers (e.g., CoqHammer) further improves AutoRocq.